OpenStudy (anonymous):

write the standard form equation of an ellipse whose vertices on the major axis are (-2,2) and (8,2) and a focus at (7,2) PLASE I NEED HELP

OpenStudy (amistre64):

we need to find the center point of the ellipse. can you do that?

OpenStudy (anonymous):

i only know that y=2

OpenStudy (anonymous):

Standard form: (x-h)^2/a^2 + (y-k)^2/b^2=1

OpenStudy (amistre64):

what is the middle between -2 and 8 ?

OpenStudy (anonymous):

4

OpenStudy (amistre64):

lets pretend it 5 :)

OpenStudy (anonymous):

ok so it is 5 lol

OpenStudy (amistre64):

wait, add and divide ...

OpenStudy (amistre64):

8-2 = 6 ... /2 = 3 lets go with 3

OpenStudy (anonymous):

ok :)

OpenStudy (amistre64):

so our center is (3,2) we can use this in the general setup that katra provided right? how would that look? and by the way, are there any options to pick from?

OpenStudy (anonymous):

no options and ok

OpenStudy (anonymous):

(x+3)^2 + (y-2)^2

OpenStudy (amistre64):

-3 nd -2 are better (x-3)^2 + (y-2)^2 -------- + ------- = 1 a^2 b^2

OpenStudy (anonymous):

why -3?

OpenStudy (amistre64):

becasue when x=3, we want that part to zero out

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

c is the distance from center to focus a is the distance from center to vertex ive got the picture bad, ill have to edit it

OpenStudy (anonymous):

ok

OpenStudy (amistre64):



OpenStudy (amistre64):

now can you use the pythagorean thrm setup to solve for a and b?

OpenStudy (anonymous):

im confused lol

OpenStudy (amistre64):

confused on what?

OpenStudy (anonymous):

how do i find the pythag if i dont know the letters

OpenStudy (amistre64):

thats why i drew the picture ... and deifned that parts

OpenStudy (anonymous):

so b is 3

OpenStudy (amistre64):

a is equal to the distance from center to vertex, what is a? c is equal to the distance from center to focus ... b is 3 yes

OpenStudy (anonymous):

c is 7

OpenStudy (amistre64):

c is 4, but immaterial since we only really need to determine a and b

OpenStudy (amistre64):

a = the distance from 3 to 8. so 5

OpenStudy (anonymous):

oh c is 5

OpenStudy (amistre64):

the confusion might be that the name of the parts of the ellipse are not quite in the same order as the abc parts in the pythag thrm

OpenStudy (amistre64):



OpenStudy (anonymous):

yes because the answer is (x-3)^2/25 + (y-2)^2/9=1

OpenStudy (amistre64):

(x-3)^2 + (y-2)^2 -------- + ------- = 1 a^2 b^2 since a=5, and b=3 ... fill them in

OpenStudy (anonymous):

25 and 9